Commercially available hydrochloric acid contains 38% HCl by mass.
Find out the volume of concentrated hydrochloric acid required to make 1litre of 0.10M HCl.(density is 1.19g/ml)

Grade:11

1 Answers

VISHAL GAUTAM
10 years ago
8 ml will be required. Fro this find the weight of hcl required . Then from this find the weight of commercial hcl required and then find the volume from the given density. Mass of hcl req=3.65gm Mass of commercial Hcl req=365/38 gm Volume req=(365/38)/1.19